# Graph Network Simulator MGN MeshGraphNet

### v1

@@ Appendix A Broader Impact
The ML-based simulator,  Rolling Diffusion-Batched Inference Network (ROBIN), offers significant advantages for computational modeling and simulation. This is achieved by significantly reducing computational costs while producing accurate simulations. This enables engineers to iterate through significantly more design variations or to quickly evaluate numerous scenarios using the fast model. However, like all powerful computational tools, there is a risk of misuse, for instance, in weapons development or unsustainable resource exploitation.
@@ Appendix B Datasets
Table 1 provides on overview over the considered datasets in this work.
Table 1: Comparison of the three datasets considered in this work. The column Nonlinearity distinguishes three different types: geometry (Geo), material (Mat), and boundary conditions (BC).
Datasets Dynamic Nonlinearity Simulator Mesh Type Steps T Dim
BendingBeam Quasi-Static Geo Scikit-fem Triangles 400 2D
ImpactPlate Dynamic Geo, BC ANSYS Triangles 52 2D
DeformingPlate Quasi-Static Geo, BC, Mat COMSOL Tetrahedrons 400 3D
BendingBeam. The dataset BendingBeam considers the bending of beam parts due to external forces. Bending is one of the most basic deformation modes of parts in structural mechanics. The dataset is designed as a diagnostic benchmark for neural PDE solvers, addressing various potential bottlenecks. The force and handle boundary conditions are very local, only beeing defined on a small subset of mesh nodes (cf. Figure 8). However, the resulting deformations affect all nodes.
Figure 8: Visualization of the different node types on the BendingBeam experiment. Blue nodes handle boundaries with fixed node positions. The red color indicates nodes where the force boundary condition is applied. Those locally applied forces in combination with the global, geometry-dependent part stiffness mainly determine the global deformation.
Hence, the neural PDE must be very good at propagating local information to all nodes of the mesh. Next, the dataset considered beams with large aspect ratios. This results in large graph diameters, which represent the shortest path between the most distant nodes. The mesh resolution is increased at thin walls, which additionally increases the graph diameter. The local boundary conditions has to be transmitted across a large number of nodes, which challenges the ability to propagate messages globally. The geometry and especially the thin locations of the geometry strongly influence the global bending stiffness and deformation of the part. Overall, the model has to output accurate solutions across various spatial frequencies.
The solutions are created with scikit-fem [61], iteratively solved using Netwon-Raphson until the residual dropped below a tolerance of $10^{-8}$. Each simulation is solved for a total number of $400$ time steps. We create a total number of $1000$ simulations for training, $100$ for validation and $100$ for testing.
@@ Appendix C Setup
Hardware and Compute. We train all models on a single NVIDIA A100 GPU with a maximum training time of $48$ hours, while most models required approximately $40$ hours. In total, we trained on $3$ datasets $9$ models each on $5$ seeds: ROBINMGNHCMT, as well as $4$ ablations and $2$ variants of ROBIN. That amounts to $40~{}\text{hours}\times 5\times 3\times 8=5400~{}\text{hours}$ training time. We required a comparable amount of time for development and hyperparameter tuning. In addition, we run inference experiments to measure the inference speed of $6$ model variants each trained on $5$ different seeds. We run the inference experiments on $3$ datasets with approximately a mean of $1$ hour per experiment. In total we obtain a runtime of $1~{}\text{hour}\times 5\times 3\times 6=90~{}\text{hours}$ for the inference experiments.
Training. We implement ROBIN in Pytorch [62] and train it with ADAM [63]. We use an exponential learning rate decay, which decrease the learning rate from $41e-$ 4 to $61e-$ 6 over the training time, including $1000$ linear increasing warm up steps. We clip gradients such that their $L_{2}$-norm doesn’t exceed $1$. We train ROBIN in BendingBeam with $9$M samples and in ImpactPlate with $6$M samples both with a batch size of $16$, resulting in $562\text{,}500$ and $375\text{,}000$ training iterations. In DeformingPlate we reduce the batch size to $12$ and train for $300\text{,}000$ iterations with $3.6$M samples.
Features. Table 2 provides an overview over the used input and ouput features for ROBIN. In addition to the default features, we extend the node embeddings of BendingBeam with the force residual $\Delta f_{\text{BC}}$, which is defined by the boundary condition. In ImpactPlate, we add the density $\rho_{i}$ and the Young’s modulus $Y_{i}$ as node features. In DeformingPlate, we add the scripted displacement residual $\Delta\mathbf{u}_{\text{BC}}$ of the actuator. We normalize all input features based on the training dataset, setting them to have a zero mean and unit variance. We add a small amount of training noise [4, 5] of $10^{-5}~{}\mathbf{\sigma}_{\mathbf{x}}$ to the node positions $\mathbf{x}_{i}^{t}$, where we scale the noise level with the standard deviation of the features $\mathbf{\sigma}_{\mathbf{x}}$. For ImpactPlate we noise the input history $\Delta\mathbf{u}^{t-1}_{i,0}=\mathbf{x}^{t}_{i}-\mathbf{x}^{t-1}_{i}$ with $10^{-3}~{}\sigma_{\mathbf{x}}$ to prevent overfitting on the history.
@@
BendingBeam $\mathbf{n}_{i}$, $\Delta\mathbf{u}^{t}_{i,k}$, $\Delta f_{\text{BC}}$ $\mathbf{n}_{i}$ ${\mathbf{x}}^{t}_{ij}$,$|{\mathbf{x}}^{t}_{ij}|$, ${\mathbf{x}}^{0}_{ij}$,$|{\mathbf{x}}^{0}_{ij}|$ ${\mathbf{x}}^{t}_{ij}$,$|{\mathbf{x}}^{t}_{ij}|$ ${\mathbf{x}}^{t}_{ij}$,$|{\mathbf{x}}^{t}_{ij}|$, ${\mathbf{x}}^{0}_{ij}$,$|{\mathbf{x}}^{0}_{ij}|$ ${\mathbf{v}}_{i,\theta}(\Delta\mathbf{u}^{t}_{i,k})$
ImpactPlate $\mathbf{n}_{i}$, $\Delta\mathbf{u}^{t}_{i,k}$, $\Delta\mathbf{u}^{t-1}_{i,0}$, $\rho_{i}$, $Y_{i}$ $\mathbf{n}_{i}$ ${\mathbf{x}}^{t}_{ij}$,$|{\mathbf{x}}^{t}_{ij}|$, ${\mathbf{x}}^{0}_{ij}$,$|{\mathbf{x}}^{0}_{ij}|$ ${\mathbf{x}}^{t}_{ij}$,$|{\mathbf{x}}^{t}_{ij}|$ ${\mathbf{x}}^{t}_{ij}$,$|{\mathbf{x}}^{t}_{ij}|$, ${\mathbf{x}}^{0}_{ij}$,$|{\mathbf{x}}^{0}_{ij}|$ ${\mathbf{v}}_{i,\theta}(\Delta\mathbf{u}^{t}_{i,k})$
DeformingPlate $\mathbf{n}_{i}$, $\Delta\mathbf{u}^{t}_{i,k}$, $\Delta\mathbf{u}_{\text{BC}}$ $\mathbf{n}_{i}$ ${\mathbf{x}}^{t}_{ij}$,$|{\mathbf{x}}^{t}_{ij}|$, ${\mathbf{x}}^{0}_{ij}$,$|{\mathbf{x}}^{0}_{ij}|$ ${\mathbf{x}}^{t}_{ij}$,$|{\mathbf{x}}^{t}_{ij}|$ ${\mathbf{x}}^{t}_{ij}$,$|{\mathbf{x}}^{t}_{ij}|$, ${\mathbf{x}}^{0}_{ij}$,$|{\mathbf{x}}^{0}_{ij}|$ ${\mathbf{v}}_{i,\theta}(\Delta\mathbf{u}^{t}_{i,k})$
Hierarchical Graph. Since the relative motion of the components in the considered experiments is not too large, we define the contact edges based on the initial mesh configuration and keep them constant to maintain a constant graph. In DeforminPlate we set the contact radius to $R=0.1$, connecting actuator nodes with plate nodes. In ImpactPlate we connect ball nodes and plate nodes with a radius of $R=1.2$. In all three experiments, we create $L=2$ coarse layers to obtain $3$ mesh levels.
Algebraic-hierarchical Message Passing Networks. We use $3$ Pre- and $3$ Post-processing layers, $2$ Up- and $2$ Downsampling layers and $5$ Solving layers, which yields a total number of $15$ learnable layers. We add a layer norm before each MLP and use two linear layers, a hidden size of $128$ and a  Sigmoid Linear Unit (SiLU[64] activation function. A max aggregation is used in all message passing layer.
Denoising Diffusion Probalistic Models. We use $K=20$ denoising steps and a denoising stride of $m=5$ for ROBIN by default. The $\beta$ variances of the DDPM scheduler are geometrically spaced for training and inference, starting from a minimum noise variance of $41e-$ 4 (for $\beta_{1}$) and going up to $1.0$ (for $\beta_{K}$).
Metrics. To compare the rollout accuracy, we follow [22] and define the Root Mean Squared Euclidean distance error $\text{RMSE}=\sqrt{1/N_{i}\sum_{i=1}^{N_{i}}\sum_{j=1}^{N_{j}}(\tilde{u}_{ij}-u_{ij})^{2}}$ between the prediction $\tilde{u}_{ij}$ and the ground truth ${u}_{ij}$ both with $N_{i}$ nodes and $N_{j}$ features. We then calculate the mean over all time steps, the mean over the dataset and finally the mean $\mu$ and standard deviation $\sigma$ over the $5$ seeds.
@@ Appendix D Baselines, Ablations and Variants
Baselines. We use the official TensorFlow [65] implementation of the authors for the baselines HCMT [^2] https://github.com/yuyudeep/hcmt/tree/main [22] and MGN [^3] https://github.com/google-deepmind/deepmind-research/tree/master/meshgraphnets [5]. We use ADAM [63] for training HCMT and MGN with an exponential leanring rate deacy form $41e-$ 4 to $61e-$ 6, a batch size of $1$ and a hidden size of $128$. We use on all datasets a total number of $15$ message passing steps for MGN, as well as a total of $15$  Hierarchical Mesh Transformer (HMT) and  Contact Mesh Transformer (CMT) layers for HCMT.
@@
On BendingBeam, we train HCMT for $4$M training iterations, use a contact radius of $R=0.01$ and a training noise [4, 5] of $0.001$. We maximize the receptive field and set the number of mesh levels to $5$, the maximum at which at least five nodes remain available across all meshes in the dataset, required for Delaunay remeshing [66]. This results in $9$ HMTs. Since BendingBeam is a contact-free task, we replaced the dual-branch CMT by $6$ single-branch Mesh Transformer layers, that only attends to mesh edges instead of mesh and contact edges. We use the same architecture and hyperparameter for HCMT on ImpactPlate and DeformingPlate as proposed by the authors [22], and train it for $2$M steps and $3$M steps, respectively.
@@
We follow the authors implementation and add world edges to the mesh graph of MGN instead of contact edges to increase the receptive field of the non-hierarchical architecture. MGN is trained for $3$M iterations on BendingBeam and uses a training noise of $0.001$ with a world edge radius of $R=0.13$. On ImpactPlate we train MGN for $3$M steps and use a world edge radius of $R=0.03$ and a training noise of $0.003$. We train MGN on DeformingPlate for $1.5$M steps and use the authors’ proposed settings [5]. To prevent out-of-memory errors in edge cases on DeformingPlate, we restrict the number of world edges to $200\text{,}000$ by selecting those with the smallest node distances.
@@
Ablations. For the No hierarchy ablation we use the fine mesh graph $\mathcal{G}^{0}$ instead of the hierarchical graph $\mathcal{G}^{0:L}$, replace our AMPN by a single Intra-MP-Stack with $15$ learnable message passing steps and remove the positional level encoding. In addition, we follow MGN and replace contact edges by world edges to increase the receptive field. To stay within the training budget, we reduce the number of trainings samples to 1.21.2{1.2}1.2M for BendingBeam and to 88{8}8M for ImpactPlate. For DeformingPlate we reduce the batch size to $1$ and the training samples to 0.60.6{0.6}0.6M and also restrict the number of world edges to $200\text{,}000$ as for MGN. The No diffusion ablation trains the AMPN with an MSE loss to predict directly the displacement residual $\Delta u^{t}$ and use an one step autoregressive rollout, such as HCMT and MGN. We use the same training noise settings as the baselines to stabilize the rollouts. The No shared layer ablation uses a total number of $15$ non-shared learnable message passing layers, that are distributed as follows: $1$ Pre-Processing and $1$ Post-Processing layer per level, $1$ Up- and $1$ Downsampling layer between each level, and $5$ Solving layers. The faster predictions allows us to increase the number of training samples to $1${11}11M for BendingBeam, to 88{8}8M for ImpactPlate, and to 4.64.6{4.6}4.6M for DeformingPlate.

### v2

@@ Appendix A Broader Impact
The ML-based simulator,  Rolling Diffusion-Batched Inference Network (ROBIN), offers significant advantages for computational modeling and simulation. This is achieved by reducing computational costs while maintaining accuracy. This enables engineers to iterate through significantly more design variations or to quickly evaluate numerous scenarios using the fast model. However, like all powerful computational tools, there is a risk of misuse, for instance, in weapons development or unsustainable resource exploitation.
@@ Appendix B Datasets
Table˜1 provides an overview of the considered datasets in this work.
Table 1: Comparison of the datasets BendingBeamBendingBeamLargeImpactPlate [22] and DeformingPlate [5] considered in this work. The column Nonlinearity distinguishes three different types: geometry (Geo), material (Mat), and boundary conditions (BC).
Datasets Dynamic Nonlinearity avg. # Nodes Mesh Type Steps T Dim
BendingBeam Quasi-Static Geo 744 Triangles 400 2D
BendingBeamLarge Quasi-Static Geo 8897 Triangles 100 2D
ImpactPlate Dynamic Geo, BC 2208 Triangles 52 2D
DeformingPlate Quasi-Static Geo, BC, Mat 1271 Tetrahedrons 400 3D
BendingBeam. This dataset considers the bending of beam parts due to external forces. Bending is one of the most basic deformation modes of parts in structural mechanics. The dataset is designed as a diagnostic benchmark for neural PDE solvers, addressing various potential bottlenecks. The force and handle boundary conditions are very local, only being defined on a small subset of mesh nodes (cf. Figure˜9). However, the resulting deformations affect all nodes.
Figure 9: Visualization of the different node types on the BendingBeam experiment. Blue nodes mark handle boundaries with fixed node positions. The red color indicates nodes where the force boundary condition is applied. Those locally applied forces in combination with the global, geometry-dependent part stiffness mainly determine the global deformation.
Hence, the neural PDE must effectively propagate local information to all nodes of the mesh. Next, the dataset considers beams with large aspect ratios. This results in large graph diameters, which represent the shortest path between the most distant nodes. The mesh resolution is increased at thin walls, which additionally increases the graph diameter. The local boundary conditions have to be transmitted across a large number of nodes, which challenges the ability to propagate messages globally. The geometry and especially the thin locations of the geometry strongly influence the global bending stiffness and deformation of the part. Overall, the model has to output accurate solutions across various spatial frequencies.
The solutions are created with scikit-fem [61], iteratively solved using Newton-Raphson until the residual fell below a tolerance of $10^{-8}$. Each simulation is solved for a total number of $400$ time steps. We create a total number of $1000$ simulations for training, $100$ for validation and $100$ for testing.
To evaluate the generalization capability of ROBIN, we create an additional dataset variant BendingBeamLarge, containing meshes more than ten times larger than those in  BendingBeam, with an average of $8897$ nodes. One simulation in BendingBeamLarge contains $100$ time steps.
@@ Appendix C Setup
Hardware and Compute. We train all models on a single NVIDIA A100 GPU with a maximum training time of $48$ hours, while most models required approximately $40$ hours. In total, we trained $11$ models across $3$ datasets, each on $5$ seeds: ROBINMGNHCMTBSMS, as well as $7$ ablations of ROBIN. That amounts to $40~\text{hours}\times 5\times 3\times 11=6600~\text{hours}$ training time. Furthermore, we trained two models on BendingBeamLarge: a pre-trained ROBIN model and a model from scratch, each on $5$ seeds for approximately $40$ hours, yielding a total duration of $40~\text{hours}\times 5\times 2=400~\text{hours}$. For each training, we required a comparable amount of time for development and hyperparameter tuning. Additionally, we conducted inference experiments to measure the inference speed for $4$ models (ROBINMGNHCMT, and BSMS) and $7$ ROBIN variants across $5$ different seeds. We run inference experiments on $3$ datasets, with each experiment taking about $1$ hour on average. In total we obtain a runtime of $1~\text{hour}\times 5\times 3\times 11=165~\text{hours}$ for the inference experiments.
Training. We implement ROBIN in PyTorch [62] and train it with ADAM [63]. We use an exponential learning rate decay, which decreases the learning rate from 1e41e-4 to 1e61e-6 over the training time, including $1000$ linearly increasing warm‑up steps. We clip gradients such that their $L_{2}$-norm doesn’t exceed $1$. We train ROBIN in BendingBeam with 99M samples and in ImpactPlate with 66M samples both with a batch size of $16$, resulting in $562\text{,}500$ and $375\text{,}000$ training iterations. In DeformingPlate we reduce the batch size to $12$ and train for $300\text{,}000$ iterations with 3.63.6M samples.
Features. Table˜2 provides an overview of the used input and output features for ROBIN. In addition to the default features, we extend the node embeddings of BendingBeam with the force residual $\Delta f_{\text{BC}}$, which is defined by the boundary condition. In ImpactPlate, we add the density $\rho_{i}$ and the Young’s modulus $Y_{i}$ as node features. In DeformingPlate, we add the scripted displacement residual $\Delta\mathbf{u}_{\text{BC}}$ of the actuator. We normalize all input features based on the training dataset, setting them to have a zero mean and unit variance. We add a small amount of training noise [4, 5] of $10^{-5}~\mathbf{\sigma}_{\mathbf{x}}$ to the node positions $\mathbf{x}_{i}^{t}$, where we scale the noise level with the standard deviation of the features $\mathbf{\sigma}_{\mathbf{x}}$. For ImpactPlate we noise the input history $\Delta\mathbf{u}^{t-1}_{i,0}=\mathbf{x}^{t}_{i}-\mathbf{x}^{t-1}_{i}$ with $10^{-3}~\sigma_{\mathbf{x}}$ to prevent overfitting on the history.
@@
BendingBeam $\mathbf{n}_{i}$, $\Delta\mathbf{u}^{t}_{i,k}$, $\Delta f_{\text{BC}}$ $\mathbf{n}_{i}$ ${\mathbf{x}}^{t}_{ij}$$|{\mathbf{x}}^{t}_{ij}|$, ${\mathbf{x}}^{0}_{ij}$$|{\mathbf{x}}^{0}_{ij}|$ ${\mathbf{x}}^{t}_{ij}$$|{\mathbf{x}}^{t}_{ij}|$ ${\mathbf{x}}^{t}_{ij}$$|{\mathbf{x}}^{t}_{ij}|$, ${\mathbf{x}}^{0}_{ij}$$|{\mathbf{x}}^{0}_{ij}|$ ${\mathbf{v}}_{i,\theta}(\Delta\mathbf{u}^{t}_{i,k})$
ImpactPlate $\mathbf{n}_{i}$, $\Delta\mathbf{u}^{t}_{i,k}$, $\Delta\mathbf{u}^{t-1}_{i,0}$, $\rho_{i}$, $Y_{i}$ $\mathbf{n}_{i}$ ${\mathbf{x}}^{t}_{ij}$$|{\mathbf{x}}^{t}_{ij}|$, ${\mathbf{x}}^{0}_{ij}$$|{\mathbf{x}}^{0}_{ij}|$ ${\mathbf{x}}^{t}_{ij}$$|{\mathbf{x}}^{t}_{ij}|$ ${\mathbf{x}}^{t}_{ij}$$|{\mathbf{x}}^{t}_{ij}|$, ${\mathbf{x}}^{0}_{ij}$$|{\mathbf{x}}^{0}_{ij}|$ ${\mathbf{v}}_{i,\theta}(\Delta\mathbf{u}^{t}_{i,k})$
DeformingPlate $\mathbf{n}_{i}$, $\Delta\mathbf{u}^{t}_{i,k}$, $\Delta\mathbf{u}_{\text{BC}}$ $\mathbf{n}_{i}$ ${\mathbf{x}}^{t}_{ij}$$|{\mathbf{x}}^{t}_{ij}|$, ${\mathbf{x}}^{0}_{ij}$$|{\mathbf{x}}^{0}_{ij}|$ ${\mathbf{x}}^{t}_{ij}$$|{\mathbf{x}}^{t}_{ij}|$ ${\mathbf{x}}^{t}_{ij}$$|{\mathbf{x}}^{t}_{ij}|$, ${\mathbf{x}}^{0}_{ij}$$|{\mathbf{x}}^{0}_{ij}|$ ${\mathbf{v}}_{i,\theta}(\Delta\mathbf{u}^{t}_{i,k})$
Hierarchical Graph. Since the relative motion of the components in the considered experiments is not too large, we define the contact edges based on the initial mesh configuration and keep them fixed to maintain a constant graph. In DeformingPlate we set the contact radius to $R=0.1$, connecting actuator nodes with plate nodes. In ImpactPlate we connect ball nodes and plate nodes with a radius of $R=1.2$. In all three experiments, we create $L=2$ coarse layers to obtain $3$ mesh levels.
Algebraic-hierarchical Message Passing Networks. We use $3$ Pre- and $3$ Post-processing layers, $2$ Up- and $2$ Downsampling layers and $5$ Solving layers, which yields a total number of $15$ learnable layers. We add a layer norm before each MLP and use two linear layers, a hidden size of $128$ and a  Sigmoid Linear Unit (SiLU[64] activation function. A max aggregation is used in all message passing layers.
Denoising Diffusion Probabilistic Models. We use $K=20$ denoising steps and a denoising stride of $m=5$ for ROBIN by default. The $\beta$ variances of the DDPM scheduler are geometrically spaced for training and inference, starting from a minimum noise variance of 1e41e-4 (for $\beta_{1}$) and going up to $1.0$ (for $\beta_{K}$).
Metrics. To compare the rollout accuracy, we follow [22] and define the Root Mean Squared Euclidean distance error $\text{RMSE}=\sqrt{1/(N_{i}N_{j})\sum_{i=1}^{N_{i}}\sum_{j=1}^{N_{j}}(\tilde{u}_{ij}-u_{ij})^{2}}$, where the prediction $\tilde{u}_{ij}$ and the ground truth ${u}_{ij}$ have $N_{i}$ nodes and $N_{j}$ features. We then calculate the mean over all time steps, the mean over the dataset and finally the mean $\mu$ and standard deviation $\sigma$ over the $5$ seeds.
@@ Appendix D Baselines, Ablations and Variants
Baselines. We use the official TensorFlow [65] implementation of the authors for the baselines HCMT [^2] https://github.com/yuyudeep/hcmt/tree/main [22] and MGN [^3] https://github.com/google-deepmind/deepmind-research/tree/master/meshgraphnets [5]. We use ADAM [63] for training HCMT and MGN with an exponential learning rate decay from 1e41e{-}4 to 1e61e{-}6, a batch size of $1$ and a hidden size of $128$. We use $15$ message‑passing steps for MGN on all datasets, as well as a total of $15$  Hierarchical Mesh Transformer (HMT) and  Contact Mesh Transformer (CMT) layers for HCMT.
@@
On BendingBeam, we train HCMT for 44M training iterations with a training noise [4, 5] of $0.001$. We maximize the receptive field and set the number of mesh levels to $5$, the maximum at which at least five nodes remain available across all meshes in the dataset, required for Delaunay remeshing [66]. This results in $9$ HMTs. Since BendingBeam is a contact-free task, we replaced the dual-branch CMT by $6$ single-branch Mesh Transformer layers that only attend to mesh edges instead of mesh and contact edges. We use the same architecture and hyperparameter for HCMT on ImpactPlate and DeformingPlate as proposed by the authors [22], and train it for 33M steps and 22M steps, respectively.
@@
We follow the authors’ implementation and add world edges to the mesh graph of MGN instead of contact edges to increase the receptive field of the non-hierarchical architecture. MGN is trained for 33M iterations on BendingBeam and uses a training noise of $0.001$ with a world edge radius of $R=0.13$. On ImpactPlate we train MGN for 33M steps and use a world edge radius of $R=0.03$ and a training noise of $0.003$. We train MGN on DeformingPlate for 1.51.5M steps and use the authors’ proposed settings [5]. To prevent out-of-memory errors in edge cases on DeformingPlate, we restrict the number of world edges to $200\text{,}000$ by selecting those with the smallest node distances.
For BSMS [48], we use the official PyTorch [62] implementation [^4] https://github.com/Eydcao/BSMS-GNN/tree/main of the authors. We follow the authors and use the maximum number of hierarchy levels possible to maximize the receptive field. We train BSMS on BendingBeam with a batch size of $12$ and 66M training samples, i.e., 500500K training iterations. We use a training noise of $0.001$ and $5$ mesh levels. On ImpactPlate we train BSMS with a batch size of $8$ and 55M samples (625625K training iterations), a training noise of $0.003$, a contact radius of $R=0.4$, and $7$ hierarchy levels. We train BSMS on DeformingPlate with a batch size of $8$ and 66M samples (750750K training iterations), a training noise of $0.003$, a contact radius of $R=0.03$, and $6$ hierarchy levels.
@@
Ablations. The 10 diffusion steps and 5 diffusion steps ablation use the same settings as ROBIN, despite the reduced number of diffusion steps $K$. For the w/o hierarchy ablation we use the fine mesh graph $\mathcal{G}^{0}$ instead of the hierarchical graph $\mathcal{G}^{0:L}$, replace our AMPN by a single Intra-MP-Stack with $15$ learnable message passing steps and remove the positional level encoding. In addition, we follow MGN and replace contact edges with world edges to increase the receptive field. To stay within the training budget, we reduce the number of training samples to 1.2{1.2}M for BendingBeam and to 8{8}M for ImpactPlate. For DeformingPlate we reduce the batch size to $1$ and the training samples to 0.6{0.6}M and also restrict the number of world edges to $200\text{,}000$ as for MGN. The w/o diffusion ablation trains the AMPN with an MSE loss to predict directly the displacement residual $\Delta\mathbf{u}^{t}_{i,0}$ and uses a one‑step autoregressive rollout, such as HCMTMGN, and BSMS. We use the same training noise settings as the baselines to stabilize the rollouts. The w/o shared layer ablation uses a total number of $15$ non-shared learnable message passing layers distributed as follows: $1$ Pre-Processing and $1$ Post-Processing layer per level, $1$ Up- and $1$ Downsampling layer between each level, and $5$ Solving layers. The faster predictions allow an increase in the number of training samples to 11{11}M for BendingBeam, to 8{8}M for ImpactPlate, and to 4.6{4.6}M for DeformingPlate. For the HCMT model ablation, we replace the AMPNs with HCMTs. More specifically, we use the same mesh hierarchy and the same model architecture as HCMTs. Everything else remains the same in ROBIN, including DDPMs and ROBI.